OpenStudy (anonymous):

Find the slope and the y- intercept. 1. 3x-4y=24

OpenStudy (zzr0ck3r):

solve for y -4y = 24 - 3x y = -6+(3/4)x y = (3/4)x-6 can you answer now?

OpenStudy (zzr0ck3r):

y = mx +b where mm is slope and b is the y intercept

OpenStudy (anonymous):

slope of equation ax+by=c is given by m=-a/b so slope of given line is m=-3/-4=3/4 for y intercept put x=0 y=24/-4=-6